hannahthespammer:

Leena consumes 400 calories at breakfast and 350 calories at lunch. She consumes StartFraction 2 Over 3 EndFraction. of her daily calories at dinner. If x represents the calories consumed at dinner, which statements describe the situation? Check all that apply.

Mercury:

statements (please post these next time) a. Leena consumed 1,500 calories at dinner. b. The equation 2/3(x + 400 + 350) = x can be used to model the situation. c. Leena consumed 500 calories at dinner. d. The equation 2/3(x) = x(400 + 300) can be used to model the situation. e. Leena consumed 1,000 calories at dinner. f. The equation 2/3x(400 + 300) = x can be used to model the situation.

Mercury:

daily calories = breakfast + lunch + dinner we also know that x = dinner calories "She consumes 2/3 of her daily calories at dinner." --> so (2/3)(daily calories) = dinner calories write the appropriate equation, then solve for x. out of the six statements, you should have two selected (one for the equation and another for the solution)